Hello,
My 8th grader is stuck on problems involving systems of equations and inequality, or substition. Can you please explain the following problem?
How to find X and Y?
Y=x+3
Y=2x+4

Answer by stanbon(48569)

Y=x+3
Y=2x+4
------------------------
Substitute to solve for "x":
x+3 = 2x+4
Subtract x from both sides.
Subtract 4 from both sides.
x = -1
--------------
Substitute to solve for "y":
y = -1+3 = 2
----------------
Solution: (-1,2)
====================
Cheers,
Stan H.